# EasyRec: Simple yet Effective Language Models for Recommendation

### v3

@@ A.2. Datasets and User/Item Profiles
In this section, we provide detailed information on the datasets, as well as the processes for profile generation and diversification, including instructions, examples, and associated costs.
A.2.1. Details of Dataset
@@
We utilize datasets from Amazon review data (Ni et al., 2019) across six categories to form the training data: Arts, Crafts and Sewing (Arts), Movies and TV (Movies), Video Games (Games), Home and Kitchen (Home), Electronics (Electronics), and Tools and Home Improvement (Tools). For the test datasets, we use one domain, Sports and Outdoors (Sports), from the Amazon review data, along with two cross-platform datasets: Steam and Yelp, for comprehensive evaluation. For the datasets from the Amazon platform, we first filter the data to include only those with a rating score greater than 3 and apply a 10-core filtering to densify the dataset. Subsequently, for each category, we split the interactions into training, validation, and test splits in a ratio of 8:1:1. In contrast, for the Steam and Yelp datasets, we directly use the data processed in previous work (Ren et al., 2024a), which maintains a split ratio of 3:1:1.
A.2.2. Details of Profile Generation
@@
After data processing, each dataset contains a split of training interactions. We use these interactions to generate user and item profiles following the paradigm described in Section 3.1, as this requires user-item interaction information. For the Steam and Yelp datasets, we directly use the provided profiles, which adhere to the same generation protocol. It is important to note that, for each dataset, the profiles are generated exclusively based on the training interactions. This approach ensures that the validation and test interactions are reserved for evaluation purposes, preventing data leakage and allowing for a more accurate assessment of the model’s generalization performance on unseen recommendation data.
The instruction prompts for generation that we utilized reference previous research (Ren et al., 2024a) to conduct efficient user and item profiling. The profiling process adopts an item-to-user paradigm, where we first generate item profiles in parallel using multi-thread processing, followed by the parallel generation of user profiles that incorporate collaborative information. The large language model employed is GPT-3.5-Turbo from OpenAI.
@@
Specifically, we process datasets from Amazon review data (Ni et al., 2019), which include the categories Arts, Movies, Games, Home, Electronics, Tools, and Sports. For item profile generation, each item $i$ includes a title $h_{i}$ and an original description $d_{i}$. We leverage these two pieces of information to generate the profile, as described on the left-hand side of Eq. 4. Next, for user profile generation, we uniformly sample a maximum of five interacted items from each user’s behavior history as collaborative information. We then arrange the input prompt for the large language model according to Eq. 5. All prompts and instructions for user and item profile generation are provided in the code for reference.
A.2.3. Details of Profile Diversification
@@
For reference, examples of user and item profile diversification are provided in Figure 6 and Figure 7, respectively. As illustrated in the case of user profile diversification, the profiles for the same user differ at the word level while still representing the same preferences. Such diversification significantly enhances the diversity and quality of the textual data while also increasing the overall dataset size.
Table 6. Costs for user and item profile generation and diversification, including three iterations of diversification.
Operation #Data #In. Tokens #Out. Tokens #Cost (\$)
Generation 7 169M 20M $\sim$114
Diversification 9 102M 30M $\sim$97
A.2.4. Cost of Generation and Diversification
We summarize the total number of tokens and the associated costs for utilizing the proprietary model for profile generation and diversification in Table 6. It is worth noting that the profiles for the Steam and Yelp datasets have already been generated; therefore, the number of profiled datasets and diversified datasets differs. As shown in the results, the total number of tokens required to process both profile generation and diversification is approximately 322 million, including both input and output tokens. This generally incurs a cost of around 200 dollars with GPT-3.5-Turbo API to process the entire dataset, making it an affordable option.
A.3. Details of Text-based Recommendation
Table 7. Details of compared language models.
@@ A.3.1. Baseline Models
  • BERT (Devlin et al., 2018): A landmark transformer-based model renowned for strong language understanding through bidirectional training. We use the pooled BERT output as the text embedding.
  • RoBERTa (Liu et al., 2019): An optimized BERT that employs dynamic masking and larger datasets. We use the final [CLS] token embedding.
  • BART (Lewis et al., 2019): A denoising autoencoder transformer trained on corrupted text reconstruction. We apply mean pooling on the last hidden state for the text embedding.
@@
  • SimCSE (Gao et al., 2021): A framework that leverages contrastive learning to generate high-quality sentence embeddings, enhancing the model’s ability to discern semantic similarity between sentences.
  • GTR (Ni et al., 2021): GTR is a generalizable T5-based dense retriever that improves retrieval tasks across various domains by overcoming limitations of traditional dual encoders.
  • BGE (Xiao et al., 2023): BGE is a state-of-the-art family of well-trained models for general text embedding models. We utilize the English version of this model for evaluation.
@@
  • BLaIR (Hou et al., 2024): A series of sentence embedding models for recommendation. BLaIR learns interaction-level correlations between item metadata and user feedback on this item, improving query-based item retrieval and also recommendation.
The pre-trained weights utilized for each baseline language model are listed in Table 7. For proprietary embedding models named OpenAIv3, we use the latest text-embedding-3-small and text-embedding-3-large from OpenAI as baselines.
A.3.2. Detail of Baseline Settings
Given that the experiment was conducted in a zero-shot setting, where the test data remained unseen during training for both our EasyRec and the other language model baselines, we directly utilized the original released parameters of these baselines from Hugging Face for initialization and comparison. Besides, for BGE, we added the recommended retrieval instruction in front of the user profiles, following the provided guidelines from the open-source code, as we found it offered better performance in zero-shot recommendations. Additionally, we applied normalization to the outputs from the language models to derive text embeddings and computed cosine similarity using the dot product of the user and item embeddings.

### v4

@@ Appendix F Datasets and User/Item Profiles
In this section, we provide detailed information on the processes for profile generation and diversification, including instructions, examples, and associated costs.
F.1 DetaILD of Dataset
@@
We utilize Amazon review data Ni et al. (2019) across six categories to form the training data: Arts, Crafts and Sewing (Arts), Movies and TV (Movies), Video Games (Games), Home and Kitchen (Home), Electronics (Electronics), and Tools and Home Improvement (Tools). For the test datasets, we use one domain, Sports and Outdoors (Sports), from the Amazon review data, along with two cross-platform datasets: Steam and Yelp, for comprehensive evaluation. For the datasets from the Amazon platform, we first filter the data to include only those with a rating score $\geq$ 3 and apply a 10-core filtering to densify the dataset. Subsequently, for each category, we split the interactions into training, validation, and test splits in a ratio of 8:1:1. In contrast, for the Steam and Yelp datasets, we directly use the data processed in RLMRec Ren et al. (2024), which maintains a split ratio of 3:1:1.
F.2 DetaILD of Profile Generation
@@
After the data processing described in Section 4.1.1, each dataset contains a split of training interactions. We use these interactions to generate user and item profiles according to the paradigm outlined in Section 3.1, as this requires user-item interaction information. For the Steam and Yelp datasets, we directly utilize the profiles generated by RLMRec Ren et al. (2024), which follow the protocol that leverages review information for user and item profile generation. It is important to note that the profiles for each dataset are generated exclusively based on the training interactions. This approach ensures that validation and test interactions are reserved for evaluation purposes, preventing data leakage and allowing for a more accurate assessment of the model’s generalization performance on unseen recommendation data.
@@
The profiling process adopts an item-to-user paradigm, where we first generate item profiles in parallel using multi-thread processing, followed by the parallel generation of user profiles that incorporate collaborative information. The large language model employed is GPT-3.5-Turbo from OpenAI. Specifically, in this work, we process our own datasets from Amazon review data Ni et al. (2019), which include the categories Arts, Movies, Games, Home, Electronics, Tools, and Sports. For item profile generation, each item $i$ includes a title $h_{i}$ and an original description $d_{i}$. We leverage these two pieces of information to generate the profile, as described on the right-hand side of Eq. 4. Next, for user profile generation, we uniformly sample a number of interacted items from each user’s behavior history as collaborative information. We then arrange the input prompt for the large language model according to Eq. 5. All prompts and instructions for user and item profile generation are provided in the code.
F.3 DetaILD of Profile Diversification
@@
For reference, examples of user and item profile diversification are provided in Figure 6 and Figure 7, respectively. As illustrated in the case of user profile diversification, the profiles for the same user differ at the word level while still representing the same preferences. Such diversification enhances the quality and diversity of textual data while increasing the overall dataset size.
Table 9: Costs for profile generation and diversification, including three iterations of diversification.
Operation #Data #Tokens #Cost (\$)
Generation 7 189M $\sim$114
Diversification 9 132M $\sim$97
F.4 Cost of Generation and Diversification
We summarize the total number of tokens and the costs for utilizing LLMs for profile generation and diversification in Table 9. The profiles for the Steam and Yelp datasets have already been generated in previous work Ren et al. (2024); therefore, the number of profiled datasets and diversified datasets differs. The total number of tokens required to process both profile generation and diversification is approximately 322 million, costing approximately 200 dollars with the GPT-3.5-Turbo API, making it a cost-effective choice.
F.5 Improved Recommendation Diversity with Profile Diversification
To further demonstrate the effecitveness of profile diversification, We also conduct experiments to explore whether diversification can enhance the diversity of recommendation results. We calculated Intra-List Distance (ILD) Zhang and Hurley (2008) for EasyRec-Large with and without profile diversification on the Steam and Yelp datasets. As shown in Table 10, our findings indicate that profile diversification effectively increases the diversity of the model’s recommendation results.
Table 10: Diversity improvement with profile diversification (Higher ILD indicates better diversity)
Model Steam Yelp
ILD@5 ILD@10 ILD@20 ILD@5 ILD@10 ILD@20
w/o Diversification 0.7128 0.7336 0.7573 0.5954 0.6090 0.6248
w/ Diversification 0.7268 0.7487 0.7727 0.6012 0.6136 0.6285
Appendix G Details of Text-based Recommendation
Table 11: DetaILD of compared language models.
@@ G.1 Baseline Models
  • BERT Devlin et al. (2018): A prominent transformer model known for its strong language understanding via bidirectional training. We use the pooled BERT output for text embedding.
  • RoBERTa Liu et al. (2019): An optimized BERT with dynamic masking and larger datasets. We use the final [CLS] token embedding.
  • BART Lewis et al. (2019): A denoising autoencoder transformer trained on corrupted text reconstruction. We apply mean pooling on the last hidden state for the text embedding.
@@
  • SimCSE Gao et al. (2021): A framework that leverages contrastive learning to generate high-quality sentence embeddings, enhancing the model’s ability to discern semantic similarity between sentences.
  • GTR Ni et al. (2021): GTR is a generalizable T5-based dense retriever that improves retrieval tasks across various domains by overcoming limitations of traditional dual encoders.
  • BGE Xiao et al. (2023): A state-of-the-art family of well-trained models for general text embeddings, utilizing the English version of BGE.
@@
  • BLaIR Hou et al. (2024a): A series of embedding models for recommendation. BLaIR learns correlations between item metadata and user feedback, improving item retrieval and recommendation.
The pre-trained weights utilized for each baseline language models are listed in Table 11. For proprietary embedding models OpenAIv3, we use the latest text-embedding-3-small and text-embedding-3-large from OpenAI.
G.2 Detail of Baseline Settings
Given that the experiment was conducted in a zero-shot setting, where the test data remained unseen during training for both our EasyRec and the other baselines, we directly utilized the original released parameters of these baselines from Hugging Face for comparison. Besides, for BGE, we added the recommended retrieval instruction in front of the user profiles, following the provided guidelines from the open-source code, as we found it offered better performance in zero-shot recommendations. We normalized model outputs and computed cosine similarity between user and item embeddings.